OpenStudy (gabylovesyou):

The height in meters of a projectile can be modeled by h = −4.9t2 + vt + s where t is the time (in seconds) the object has been in the air, v is the initial velocity (in meters per seconds) and s is the initial height (in meters). A soccer ball is kicked upward from the ground and flies through the air with an initial vertical velocity of 4.9 meters per second. After how many seconds does it land approximately? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

in this case, v = 4.9 and s = 0, so the equation is h = −4.9t2 + 4.9t

OpenStudy (gabylovesyou):

t = 1 ?

jimthompson5910 (jim_thompson5910):

what you need to do from here is plug in h = 0 and solve for t

jimthompson5910 (jim_thompson5910):

correct, hopefully that's not a random guess and you know how you got that?

OpenStudy (gabylovesyou):

lol yes i do know :p lmfao xDh = -4.9t² + 4.9t and when it's on the ground again, h = 0, so -4.9t² + 4.9t = 0 -4.9t (t – 1) = 0 t = 1 sec.

jimthompson5910 (jim_thompson5910):

very nice work, just checking to make sure you got it